# MGAS: Multi-Granularity Architecture Search for Trade-Off Between Model Effectiveness and Efficiency

### v3

@@ Abstract
Neural architecture search (NAS) has gained significant traction in automating the design of neural networks. To reduce the time cost, differentiable architecture search (DAS) transforms the traditional paradigm of discrete candidate sampling and evaluation into that of differentiable super-net optimization and discretization. However, existing DAS methods fail to trade off between model performance and model size. They either only conduct coarse-grained operation-level search, which results in redundant model parameters, or restrictively explore fine-grained filter-level and weight-level units with pre-defined remaining ratios, suffering from excessive pruning problem. Additionally, these methods compromise search quality to save memory during the search process. To tackle these issues, we introduce multi-granularity architecture search (MGAS), a unified framework which aims to discover both effective and efficient neural networks by comprehensively yet memory-efficiently exploring the multi-granularity search space. Specifically, we improve the existing DAS methods in two aspects. First, we balance the model unit numbers at different granularity levels with adaptive pruning. We learn discretization functions specific to each granularity level to adaptively determine the unit remaining ratio according to the evolving architecture. Second, we reduce the memory consumption without degrading the search quality using multi-stage search. We break down the super-net optimization and discretization into multiple sub-net stages, and perform progressive re-evaluation to allow for re-pruning and regrowing of previous units during subsequent stages, compensating for potential bias. Extensive experiments on CIFAR-10, CIFAR-100 and ImageNet demonstrate that MGAS outperforms other state-of-the-art methods in achieving a better trade-off between model performance and model size.
Index Terms:
Neural architecture search (NAS), pruning, deep neural networks (DNNs), image classification
footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Figure 1: Illustration of the accuracy-parameter trade-off with respect to different remaining ratios for units at different granularity levels (operation, weight and filters) on CIFAR-10 dataset. In (a), we maintain a constant operation number and manipulate the filter number and sparsity ratio, while ensuring the model parameters remain unchanged at approximately 2M. In (b), we keep the sparsity ratio fixed and modify the operation number and filter number, while maintaining the model parameters at around 2.3M. We observe that model accuracy varies significantly even when model sizes are similar. This underscores the necessity of effectively balancing the units of different granularities.
@@
Neural architecture search, the method for automatically designing effective and efficient neural networks, has gained significant popularity in various domains including computer vision [1, 2, 3] and natural language processing [1, 4]. In the early stage of NAS, architectures were heuristically sampled from a search space using reinforcement learning [5, 6] or evolutionary algorithms [7, 8], followed by individual training and performance comparison, which incurred significant computational cost. Weight-sharing techniques were later introduced to alleviate the burden by enabling the sharing of model weights among different architectures through a super-network created from the search space. By training only a single super-net, the computational resources required can be greatly reduced. Based on that, differentiable architecture search (DAS), e.g., DARTS [1], further improved the efficiency by introducing differentiable architecture parameters which can be optimized along with the super-net weights via gradient descent. This advancement allows for the direct updates of the architecture during super-net training, transforming the search process into the optimization and discretization of the super-net.
@@
In spite of the great success of DAS, the current research mainly focuses on coarse-grained search in constrained search spaces that only cover operation-level searchable units. While this limited granularity stabilizes the search process, it inevitably results in the presence of redundant parameters in the discovered architectures, which consequently affects model efficiency.
@@
To eliminate the redundancy, recent studies have emerged to expand the search space and investigate finer-grained units at the filter and weight levels. While certain studies [9, 10, 11] perform filter-level search, their methods are limited by determining the remaining filter ratio from a restricted set of options, which lacks flexibility. Moreover, they cannot be extended to weight-level search which offers a much broader range of sparsity pattern candidates. On the other hand, [12, 13] directly prune on weight-level units until the network reaches a manually-defined target sparsity ratio, aiming to discover well-performing highly-sparse architectures. However, it is worth noting that the sparsity ratio itself is a crucial factor influencing the model performance. It is also mentioned in [14] that there exists an optimal sparsity ratio for a specific neural network in terms of achieving the best performance. We further reveal that such performance discrepancies cannot be solely attributed to the differences in model sizes introduced by sparsity. Even models with similar parameter counts can exhibit significant performance variations depending on different remaining ratios of the operations, filters and weights, as shown in Fig. 1. This observation highlights the necessity of balancing the unit numbers at different granularity levels. Failure to consider the relative numbers of the multi-granularity units can lead to excessive pruning at certain levels, which degrades the model effectiveness.
@@
Additionally, DAS requires the whole super-net to be placed in the memory for optimization and discretization, leading to substantial memory usage. This issue becomes more pronounced when the finer-grained searchable units are introduced. For example, in order to facilitate the flexible filter-level search, the initial super-net must possess a larger number of filters for candidate selection, which increases the demands of the memory. Several techniques have been proposed to save memory. The most commonly applied proxy strategy involves optimizing a shallower [1, 15, 16] or slimmer [17, 18] proxy super-net during the search, followed by extension of the discovered architecture to form the final target network. However, this configuration change renders the proxy strategy unsuitable for flexible fine-grained search. Another approach is the decomposition strategy, which factorizes the super-net along the path [19, 20, 21] or the depth [22, 23, 24] for separate optimization. Nevertheless, this strategy introduces bias and compromises the search quality, as the searchable units are prematurely evaluated without considering the overall architecture. As a result, there is a necessity to propose a more general and effective memory reduction strategy applicable to flexible and complex search spaces.
To address these issues, we propose multi-granularity architecture search (MGAS), which serves as a unified framework to comprehensively explore the multi-granularity search space in a memory-efficient way, aiming to discover neural networks that are both effective and efficient. Specifically, we balance the unit numbers of different granularities through an adaptive pruning scheme. This involves learning discretization functions specific to each granularity level during the search to adaptively adjust the pruning criteria for different granularities according to the evolving architecture. Consequently, the remaining ratios of units at different granularity levels can be properly determined for models at different target sizes. To reduce the memory consumption for this search space, we propose a multi-stage search strategy. This entails decomposing the super-net optimization and discretization into sequential sub-net stages. In each stage, the size of preceding sub-nets is significantly reduced, ensuring low memory usage. However, since the preceding sub-nets are greedily optimized without considering the structure of subsequent ones, the early stages are inevitably biased like existing decomposition methods. To compensate for the bias, we further propose to progressively re-evaluate the previous units as we move to the later stages, so that the remaining units in preceding sub-nets can undergo additional pruning, and the pruned units have the opportunity to regrow based on the subsequent sub-nets.
@@
    To the best of our knowledge, we conduct the first in-depth study on balancing the unit numbers at different granularity levels in NAS to trade off between model effectiveness and model efficiency.
@@
    We demonstrate the superiority of MGAS through extensive experiments on CIFAR-10, CIFAR-100 and ImageNet. Compared with baseline methods, MGAS achieves a better trade-off between model performance and model size.
@@ II-A Search Space Expansion
For the search efficiency, prevalent DAS approaches often perform the search in cell-based search spaces [25], e.g., DARTS search space. Recent works relax the constraints within these search spaces to discover novel designs.
To enable flexible connection patterns, Amended-DARTS [26] explores independent cell structures. GOLD-NAS [27] allows each node to preserve an arbitrary number of precedents and each edge to preserve an arbitrary number of operations.
Other works turn to investigate finer-grained searchable units. Single-path NAS [9] jointly searches for the kernel size and expansion ratio by encoding all the candidate convolution operations into one single ’superkernel’. Trilevel NAS [11] also defines a set of searchable expansion ratios and performs the filter-level search with parameters indicating the probability of the choices. However, both works lack flexibility in determining the filter number as they are limited to several pre-defined options. Moreover, they cannot be extended to search for the various sparsity patterns that cannot be encoded into several options at the weight level. In comparison, SuperTickets [12] iteratively prunes unimportant filters and weights to identify efficient networks and their ’lottery subnetworks’. DASS [13] also directly utilizes pruning technique, and explores operation-level and weight-level units by incorporating two sparse candidate operations to find sparsity-friendly networks. Nevertheless, both works pre-define the desired sparsity ratio and ignore the exploration of the unit distribution at different granularity levels.
In this work, we follow the connection pattern settings in GOLD-NAS due to its high flexibility, and integrate operation-level, filter-level and weight-level units to form a multi-granularity search space, targeting at discovering a network that achieves a better accuracy-parameter trade-off. We specifically focus on identifying proper remaining ratios for units of different granularities.
@@ II-B Memory Reduction Strategies
DARTS [1] and its subsequent works [15, 16, 13] apply a depth-level proxy strategy to save memory by searching for a network with repeatable cells in a shallower proxy super-net, and constructing a deeper target network with stacked cells. To fill the performance gap caused by the inconsistency between the proxy and target networks, PDARTS [28] progressively reduces the candidate operations and increases the network depth. PC-DARTS [17] applies the proxy strategy at the width level by sampling a sub-set of channels for optimization while bypassing the held out part in a shortcut. However, the extension of depth necessitates identical cell structures in the network and the extension of width requires identical sparsity patterns in the filters. Consequently, the proxy strategy is not applicable for conducting flexible fine-grained search in our multi-granularity search space.
In contrast, ProxylessNAS [19] proposes a path binarization strategy to activate only one or two paths in the super-net for optimization at each iteration. GDAS [20] utilizes the Gumbel-Max trick for the discrete sampling of the single-path architecture. To reduce the instability brought by sampling, MSG-DAS [21] extends the Gumbel sampler to optimize multiple mutually exclusive single-path architectures at each iteration. However, these methods cannot discover well-performed multi-path architectures due to their biased optimization of operations without considering counterparts on the same edge. DNA [22] factorizes the super-net along the depth into blocks and employs a pre-trained teacher model for block-wise supervision to enable separate optimization. BossNAS [23] uses self-supervised learning as an alternative to the supervised distillation. Sharing a similar decomposition idea, PNAS [24] starts with a small number of blocks, progressively removing the unpromising structures in previous blocks and extending the block number. However, the importance of previously evaluated units may vary in the complete architecture, which also introduces bias in the search process.
In this work, we aim to propose a general memory reduction strategy applicable to the multi-granularity search space yet effective enough to ensure high search quality.
@@ III-1 DARTS Search Space
The original DARTS search space consists of multiple cells, wherein each cell is a directed acyclic graph encompassing an ordered sequence of $M$ computational nodes. Each node $x_{i}$ serves as a latent representation and each directed edge $(i,j)$ denotes the information transformation from $x_{i}$ to $x_{j}$. The operations $o$ on each edge need to be determined, where the operation types include convolutions with different kernel sizes, pooling and skip connection.
@@ IV Methodology
In this section, we first describe the multi-granularity search space to be explored. Then, we introduce our proposed MGAS framework, which consists of 1) adaptive pruning to balance the remaining ratios of units at different granularity levels in the discovered networks, and 2) multi-stage search to reduce the memory usage without compromising the search quality.
Figure 3: Overview of pruning with fixed pruning rates and adaptive pruning. The dotted arrows with different colours represent the pruning process at different granularity levels. The width of the dotted arrow indicates the number of the pruned units. Existing works independently prune on different granularities with fixed pruning rates. In contrast, our adaptive pruning learn granularity-specific discretization functions to adaptively determine the pruning number at each granularity level according to the architecture evolution, so that the relative remaining numbers at different levels can be optimized for models with different sizes.
@@ IV-A Multi-Granularity Search Space
The problem of searching for the proper filter number and sparsity pattern can be viewed as that of selecting the important subset of filters and weights from the super-net. The differentiable architecture parameter $\alpha$ can be viewed as the operation-level granularity parameter that represent the operation importance. Inspired by this, we introduce differentiable filter-level granularity parameter denoted as $\beta$ to indicate the significance of each filter. For an individual weight, we directly use the value of the weight itself $\omega$ to represent its importance. Mathematically, in each convolution operation $o(x_{i})$ in the constructed super-net, there is $C_{out}^{f}=\sigma(\beta^{f})\cdot\omega^{f}\cdot C_{in}$, where $C_{in}$ is the input feature map, $f$ refers to a filter, $C_{out}^{f}$ is the feature map generated by $f$, $\omega^{f}$ denotes the weights in the filter, $\beta^{f}$ represents the granularity parameter of $f$, and $\sigma(\beta^{f})$ is the scaled granularity parameter. In this super-net, we need to optimize and discretize the scaled granularity parameters of the three levels, i.e., $\sigma(\alpha)$, $\sigma(\beta)$ and $\omega$.
@@ IV-B Adaptive Pruning
As discussed above, the primary concern in exploring the multi-granularity search space is the joint optimization and discretization of the scaled granularity parameters $\sigma(\alpha)$, $\sigma(\beta)$ and $\omega$. Previous studies [12, 13] have concentrated on improving joint optimization to accurately compare the unit importance within each granularity level. They achieve this by adaptively updating the granularity parameters of different levels, taking into account the changes in unit importance caused by the evolution of units at other levels. However, they independently discretize the granularity parameters at different levels with fixed pruning rates, failing to optimize the relative numbers of the remaining multi-level units. In this case, we focus on improving joint discretization with adaptive pruning to adaptively determine the relative numbers according to the evolving architecture. To achieve this, we need to enable the comparison of unit importance across different granularities. Given that the granularity parameters are measured on different scales, a direct numerical comparison is not feasible. Therefore, we propose the dynamic acquisition of granularity-specific pruning criteria to determine when a unit can be considered unimportant within its own granularity level along with the architecture evolution, enabling implicit comparison of unit importance across granularities.
Technically, instead of manually defining pruning rates or importance thresholds for discretization, we learn granularity-specific discretization functions $M_{\alpha}(\cdot),M_{\beta}(\cdot),M_{\omega}(\cdot)$ during the super-net optimization, as shown in Fig. 3. The objective of the discretization functions is to zero out the scaled granularity parameters $\sigma(\alpha),\sigma(\beta),\omega$ whose magnitudes are small enough. We formulate them as
@@
where $S(\cdot)$ is a binary step function, $t_{\alpha},t_{\beta},t\omega$ are trainable thresholds. To make $S(\cdot)$ differentiable, we follow [29] to adopt a long-tailed higher-order estimator.
@@
$C_{out}^{f}=M_{\beta}^{o}(\sigma(\beta^{f}))\cdot M_{\omega}^{o}(\omega^{f})\cdot C_{in}.$ (5)
By learning the discretization functions $M_{\alpha}(\cdot),M_{\beta}(\cdot),M_{\omega}(\cdot)$, we obtain tailored pruning criteria for each granularity adaptively according to the evolving architecture, so that the unit remaining ratios can be properly determined.
@@
Input: dataset $D$, pruning interval $e$, target model size $Param_{min}$, initial remaining units $U_{act}$
@@
4:     Update loss $L\leftarrow Loss(D;\alpha^{U_{act}},\beta^{U_{act}},\omega^{U_{act}})$triangleright Optimization
@@
7:     for each operation $o$ in $U_{act}$ do $\triangleright$ Discretization
8:         Update $\sigma(\alpha^{o})$ with discretization function $M_{\alpha}^{o}(\cdot)$: $\sigma(\alpha^{o})\leftarrow S(|\sigma(\alpha^{o})|-t_{\alpha}^{o})\cdot\sigma(\alpha^{o})$
9:         if $epoch\bmod e=0$ then
10:              if $\sigma(\alpha^{o})=0$ then
11:                  Update remaining units $U_{act}\leftarrow U_{act}\setminus o$
12:                  continue
13:              end if
14:         end if
15:         for each filter $f$ in $o$ do
16:              Update $\sigma(\beta^{f})$ with discretization function $M_{\beta}^{o}(\cdot)$: $\sigma(\beta^{f})\leftarrow S(|\sigma(\beta^{f})|-t_{\beta}^{o})\cdot\sigma(\beta^{f})$
17:              if $epoch\bmod e=0$ then
18:                  if $\sigma(\beta^{f})=0$ then
19:                       Update remaining units $U_{act}\leftarrow U_{act}\setminus f$
20:                       continue
21:                  end if
22:              end if
23:              for each weight $w$ in $f$ do
24:                  Update $\omega^{w}$ with discretization function $M_{\omega}^{o}(\cdot)$: $\omega^{w}\leftarrow S(|\omega^{w}|-t_{\omega}^{o})\cdot\omega^{w}$
25:                  if $epoch\bmod e=0$ then
26:                       if $\omega^{w}=0$ then
27:                           Update remaining units $U_{act}\leftarrow U_{act}\setminus w$
28:                           continue
29:                       end if
30:                  end if
31:              end for
32:         end for
33:     end for
34:end while

### v4

# Xiaoyun Liu1, Divya Saxena2, Jiannong Cao1, Yuqing Zhao1, Penghui Ruan1

@@ Abstract
Neural architecture search (NAS) has gained significant traction in automating the design of neural networks. To reduce search time, differentiable architecture search (DAS) reframes the traditional paradigm of discrete candidate sampling and evaluation into a differentiable optimization over a super-net, followed by discretization. However, most existing DAS methods primarily focus on optimizing the coarse-grained operation-level topology, while neglecting finer-grained structures such as filter-level and weight-level patterns. This limits their ability to balance model performance with model size. Additionally, many methods compromise search quality to save memory during the search process. To tackle these issues, we propose Multi-Granularity Differentiable Architecture Search (MG-DARTS), a unified framework which aims to discover both effective and efficient architectures from scratch by comprehensively yet memory-efficiently exploring a multi-granularity search space. Specifically, we improve the existing DAS methods in two aspects. First, we adaptively adjust the retention ratios of searchable units across different granularity levels through adaptive pruning, which is achieved by learning granularity-specific discretization functions along with the evolving architecture. Second, we decompose the super-net optimization and discretization into multiple stages, each operating on a sub-net, and introduce progressive re-evaluation to enable re-pruning and regrowth of previous units, thereby mitigating potential bias. Extensive experiments on CIFAR-10, CIFAR-100 and ImageNet demonstrate that MG-DARTS outperforms other state-of-the-art methods in achieving a better trade-off between model accuracy and parameter efficiency. Codes are available at https://github.com/lxy12357/MG_DARTS.
@@ I Introduction
Figure 1: Illustration of the accuracy-parameter trade-off with respect to different retention ratios for units at different granularity levels (operation, weight and filters) on CIFAR-10 dataset. In (a), we maintain a constant operation number and manipulate the filter number and sparsity ratio, while ensuring the model parameters remain unchanged at approximately 2M. In (b), we keep the sparsity ratio fixed and modify the operation number and filter number, while maintaining the model parameters at around 2.3M. We observe that model accuracy varies significantly even when model sizes are similar. This underscores the necessity of effectively balancing the units of different granularities.
@@
Neural architecture search, the method for automatically designing effective and efficient neural networks, has gained significant popularity in various domains including computer vision [1, 2], natural language processing [1, 3]. In the early stage of NAS, architectures were heuristically sampled from a search space using reinforcement learning [4, 5] or evolutionary algorithms [6, 7], followed by individual training and performance comparison, which incurred significant computational cost. Weight-sharing techniques were later introduced to alleviate the burden by enabling the sharing of model weights among different architectures through a super-network created from the search space. By training only a single super-net, the computational resources required can be greatly reduced. Based on that, differentiable architecture search (DAS), e.g., DARTS [1], further improved the efficiency by introducing differentiable architecture parameters which can be optimized along with the super-net weights via gradient descent. This advancement allows for the direct updates of the architecture during super-net training, transforming the search process into the optimization and discretization of the super-net.
@@
In spite of the great success of DAS, balancing model performance and model size remains a critical challenge in NAS [8, 9, 10]. The current research mainly focuses on coarse-grained search in constrained search spaces that only cover operation-level searchable units. While this limited granularity stabilizes the search process, it inevitably results in the presence of redundant parameters in the discovered architectures, which consequently affects model efficiency.
@@
To eliminate the redundancy, recent studies have emerged to expand the search space and jointly investigate finer-grained units at the filter and weight levels. While certain studies explore filter-level search [11, 12, 13, 14, 15] and weight-level search [16, 17], most of them overlook the coordination of retention ratios of units at different granularities. However, the retention ratio itself plays a critical role in determining model performance. For example, it is mentioned in [18] that there exists an optimal sparsity ratio for a specific neural network in terms of achieving the best performance. We further observe that such performance discrepancies cannot be solely attributed to the differences in model sizes introduced by sparsity. Even models with similar parameter counts can exhibit great performance variations depending on different retention ratios of the operations, filters and weights, as shown in Fig. 1. This underscores the importance of adaptively adjusting unit proportions across different granularity levels. Failing to consider this may result in excessive pruning at certain levels, particularly in the absence of full super-net pretraining, ultimately degrading model performance.
@@
Additionally, DAS requires the whole super-net to be placed in the memory for optimization and discretization, leading to substantial memory usage. This issue becomes more pronounced when the finer-grained searchable units are introduced. For example, in order to facilitate the flexible filter-level search, the initial super-net must possess a larger number of filters for candidate selection, which increases the demands of the memory. Several techniques have been proposed to reduce memory consumption during search. Path-level decoupling strategies [19, 20, 21, 22, 23, 24] activate only a limited subset of operations at each iteration, but often overlook inter-path interactions, limiting the discovery of high-performing multi-path architectures. Depth-level [1, 25, 26, 27, 28, 17, 29] and width-level [30, 31, 32, 33, 34, 35, 36, 37, 29] strategies simplify or decompose the network structure during search, but may introduce bias, as the searchable units are prematurely evaluated without considering the overall architecture. As a result, there is a necessity to propose a more general and effective memory reduction strategy applicable to flexible and complex search spaces.
To address these issues, we propose Multi-Granularity Differentiable Architecture Search (MG-DARTS), a unified framework that enables comprehensive and memory-efficient exploration of a multi-granularity search space, aiming to achieve a better trade-off between model performance and size. Specifically, we introduce a granularity-aware adaptive pruning mechanism that jointly prunes operations, filters, and weights within a unified framework. This is achieved by learning discretization functions specific to each granularity level, allowing the pruning criteria to adapt to the evolving architecture. As a result, MG-DARTS can flexibly determine appropriate retention ratios across different granularities to meet various target model sizes. To reduce memory consumption, we further propose a multi-stage search strategy that decomposes super-net optimization and discretization into a sequence of sub-net stages. Each stage operates on a reduced sub-net, maintaining low memory usage. However, early-stage pruning may introduce bias, as it is performed without foresight of subsequent architectural choices. To address this, we introduce progressive re-evaluation, allowing previously pruned units to be reconsidered and previously retained units to be further pruned, thereby mitigating bias across stages.
@@
    To the best of our knowledge, we conduct the first in-depth study on adaptively adjusting the retention ratio at different granularity levels in NAS to trade off between model effectiveness and model efficiency.
@@
    We demonstrate the superiority of MG-DARTS through extensive experiments on CIFAR-10, CIFAR-100 and ImageNet. Compared with baseline methods, MG-DARTS achieves a better trade-off between model performance and model size.
@@ II-A Search Space Expansion
Considering search efficiency, most DAS approaches adopt cell-based search spaces [38] such as the widely used DARTS search space. Recent works have attempted to relax the constraints within these search spaces to enable the discovery of more diverse and effective network designs.
To enable more flexible network topologies, Amended-DARTS [39] explores independent cell structures. GOLD-NAS [40] further allows each node to preserve an arbitrary number of precedents and each edge to preserve an arbitrary number of operations.
Motivated by pruning techniques, other works turn to introduce finer-grained searchable units. At the filter level, FBNetV2 [13] defines a set of candidate blocks with varying filter numbers for each layer and formulates filter number selection as a differentiable process. Single-Path NAS [11] encodes all candidate convolution operations into a single ‘superfilter’ and jointly searches for the optimal kernel size and expansion ratio by applying learnable thresholds over different subsets. TangleNAS [14] enhances this approach by introducing additional architecture parameters to better capture the importance of each subset. ReCNAS [15] proposes dynamic programming-based channel pruning to enable more fine-grained filter pattern selection. Beyond filter-level granularity, weight-level search has also been explored. For example, SuperTickets [16] iteratively prunes unimportant operations and weights to identify the efficient architectures and their ‘lottery sub-networks’ in an end-to-end way. DASS [17] incorporates two sparse candidate operations to promote sparsity-friendly network discovery.
However, most of these methods focus primarily on optimizing unit distribution within individual granularity levels, while ignoring the allocation trade-offs across different levels of granularity.
In this work, we follow GOLD-NAS to relax the constraints on network topology to enable variation in the final operation number within the architecture. Building on this flexibility, we incorporate operation-level, filter-level, and weight-level units to construct a unified multi-granularity search space. On top of this, we propose a general differentiable architecture search framework to jointly optimize the retention ratios of units across different granularities, with the goal of discovering architectures that achieve a better trade-off between accuracy and parameter efficiency.
@@ II-B Memory Reduction Strategies
Some works adopt path-level decoupling strategies, where only a subset of operations is activated at each iteration. ProxylessNAS [19] proposes a path binarization strategy to activate only one or two paths in the super-net. GDAS [20] and D2NAS [21] employ the Gumbel-Max trick for discrete path sampling. On-NAS [23] adopts an expectation-based sampling approach over both edge pairs and operations. U-DARTS [22] applies random sampling. FP-DARTS [24] partitions candidate operations into two disjoint subsets and treats each subset as a unified path. However, these approaches are limited in discovering well-performing multi-path architectures due to biased optimization that neglects the interaction among operations on the same edge.
In contrast, depth-level proxy or decomposition strategies aim to reduce search cost by operating on shallower networks. DARTS [1] and its subsequent works [25, 17] search over a shallow super-net with repeatable cells and later stack them to form a deeper target network. To fill the performance gap caused by this proxy-target mismatch, PDARTS [41], PNAS [28], and MIG-DARTS [29] progressively reduce the number of candidate operations while increasing the network depth. DNA [26] factorizes the super-net along the depth axis into blocks and applies block-wise supervision with a pre-trained teacher model. BossNAS [27] instead leverages self-supervised learning for block-wise guidance.
Similarly, width-level proxy strategies improve efficiency by reducing the number of active channels during search. PC-DARTS [30] and its subsequent works [32, 33] samples a subset of channels for operation selection while bypassing the held out part in a shortcut. EPC-DARTS [34] and ACA-PC-DARTS [35] enhance this by introducing attention modules for channel selection. $l$-DARTS [36] reparameterizes mixed operation outputs through weighted normalization to improve stability. PA-DARTS [37] and MIG-DARTS [29] further apply progressive channel scaling to alleviate the performance gap.
However, both depth-level and width-level strategies may introduce bias, as the importance of previously evaluated units may vary in the full-scale architecture.
There are also some works which directly limits the number of candidate operations [39, 40], starts from a small set [42], or explore more parameter-efficient candidate operations [43].
In this work, we propose multi-stage search, a depth-level memory reduction strategy augmented with a re-evaluation mechanism to mitigate the aforementioned bias.
@@ III-1 DARTS Search Space
The original DARTS search space consists of multiple cells, wherein each cell is a directed acyclic graph encompassing an ordered sequence of $M$ computational nodes. Each node $x_{i}$ serves as a latent representation and each directed edge $(i,j)$ denotes the information transformation from $x_{i}$ to $x_{j}$. The operations $o$ on each edge need to be determined, where the operation types include convolutions with different filter sizes, pooling and skip connection.
@@ IV Methodology
In this section, we first describe the multi-granularity search space to be explored. Then, we introduce our proposed MG-DARTS framework, which consists of 1) adaptive pruning to balance the retention ratios of units at different granularity levels in the discovered networks, and 2) multi-stage search to reduce the memory usage without compromising the search quality.
Figure 3: Overview of pruning with fixed pruning rates and adaptive pruning. The dotted arrows with different colours represent the pruning process at different granularity levels. The width of the dotted arrow indicates the number of the pruned units. Existing works independently prune on different granularities with fixed pruning rates. In contrast, our adaptive pruning learns granularity-specific discretization functions to adaptively determine the pruning ratio at each granularity level according to the architecture evolution, so that the retention ratios at different levels can be optimized for models with different sizes.
@@ IV-A Multi-Granularity Search Space
The problem of searching for the proper filter number and sparsity pattern can be viewed as that of selecting the important subset of filters and weights from the super-net. This naturally unifies the search across different granularity levels into the same subset selection paradigm used in traditional NAS. The differentiable architecture parameter $\alpha$ can be viewed as the operation-level granularity parameter that represent the operation importance. Inspired by this, we introduce differentiable filter-level granularity parameter denoted as $\beta$ to indicate the significance of each filter. For an individual weight, we directly use the magnitude of the weight itself $\omega$ to represent its importance. Mathematically, in each convolution operation $o(x_{i})$ in the constructed super-net, there is $C_{out}^{f}=\sigma(\beta^{f})\cdot\text{Conv}(C_{\text{in}},\omega_{f})$, where $C_{in}$ is the input feature map, $f$ refers to a filter, $C_{out}^{f}$ is the feature map generated by $f$, $\omega^{f}$ denotes the weights in the filter, $\beta^{f}$ represents the granularity parameter of $f$, and $\sigma(\beta^{f})$ is the scaled granularity parameter. In this super-net, we need to optimize and discretize the scaled granularity parameters of the three levels, i.e., $\sigma(\alpha)$, $\sigma(\beta)$ and $\omega$.
@@ IV-B Adaptive Pruning
As discussed above, the primary concern in exploring the multi-granularity search space is the joint optimization and discretization of the scaled granularity parameters $\sigma(\alpha)$, $\sigma(\beta)$ and $\omega$. Previous studies [16, 17] have concentrated on improving joint optimization to accurately compare the unit importance within each granularity level. They achieve this by adaptively updating the granularity parameters of different levels, taking into account the changes in unit importance caused by the evolution of units at other levels. However, they independently discretize the granularity parameters at different levels with fixed pruning rates, failing to optimize the retention ratios of multi-level units. In this case, we focus on improving joint discretization with adaptive pruning to adaptively determine the retention ratios according to the evolving architecture. To achieve this, we need to enable the comparison of unit importance across different granularities. Given that the granularity parameters are measured on different scales, a direct numerical comparison is not feasible. Therefore, we propose the dynamic acquisition of granularity-specific pruning criteria to determine when a unit can be considered unimportant within its own granularity level along with the architecture evolution, enabling implicit comparison of unit importance across granularities.
Technically, instead of manually defining pruning rates or importance thresholds for discretization, we learn granularity-specific discretization functions $M_{\alpha}(\cdot),M_{\beta}(\cdot),M_{\omega}(\cdot)$ with differentiable thresholds $t_{\alpha},t_{\beta},t_{\omega}$ during the super-net optimization, as shown in Fig. 3. The objective of the discretization functions is to zero out the scaled granularity parameters $\sigma(\alpha),\sigma(\beta),\omega$ whose magnitudes are small enough. We formulate them as
@@
where $S(\cdot)$ is a binary step function, $t_{\alpha},t_{\beta},t_{\omega}$ are trainable thresholds. To make $S(\cdot)$ differentiable, we follow [44] to adopt a long-tailed higher-order estimator.
@@
$C_{out}^{f}=M_{\beta}^{o}(\sigma(\beta^{f}))\cdot\text{Conv}(C_{in},M_{\omega}^{o}(\omega^{f})).$ (5)
By learning the discretization functions $M_{\alpha}(\cdot),M_{\beta}(\cdot),M_{\omega}(\cdot)$, we obtain tailored pruning criteria for each granularity adaptively according to the evolving architecture, so that the unit retention ratios can be properly determined.
@@
Input: dataset $D$, target model size $Param_{min}$, initial remaining units $U_{act}$
@@
4:  Update loss $L\leftarrow Loss(D;\alpha^{U_{act}},\beta^{U_{act}},\omega^{U_{act}})$ $\triangleright$ Optimization
@@
7:  for each granularity param $g$ in $\alpha,\beta,\omega$ do $\triangleright$ Discretization
8:   Update $\sigma(g)$ with discretization function $M_{g}(\cdot)$: $\sigma(g)\leftarrow S(|\sigma(g)|-t_{g})\cdot\sigma(g)$
9:   for unit $u$ in $U_{act}$ do
10:     if $\sigma(g^{u})=0$ then
11:      Update remaining units $U_{act}\leftarrow U_{act}\setminus u$
12:     end if
13:   end for
14:  end for
15:end while